# Chiral dynamics with giant systems

### v1

@@ Abstract
Figure 3: DPT occurs only when the giant system is driven by right-moving input. (a) Photon number (blue solid line, left axis) in the cavity and the second-order correlation function (red line with circles, right axis), as functions of the driving strength (b)Photon number density (blue lines, left axis) of the reflected field and the two-point correlation (red lines with circles, right axis). for different distance $\gamma d=0.63$ (solid lines), $\gamma d=63$ (dashed lines), and $\gamma d=6300$ (dotted lines). Other parameters are $U=\gamma$ and $k_{i}=10\gamma$.
In this work, we have investigated the dynamics of a driven giant system under chiral conditions, specifically when the phase difference $\phi$ between the two coupling points matches the propagation phase at the driving frequency. We focused on a giant cavity with nonlinearity $U$ that interacts solely with input light from a single direction. In the weak-driving regime, equivalent to a few-photon scattering process, we demonstrated that photons with nontrivial statistics can be deterministically generated, even in the strong dissipation regime where $U$ is small. This is due to the interference in the reflection and transmission of photons propagating between the two coupling points. Notably, when $\phi=\pm\pi/2$, this interference vanishes, resulting in Markovian dynamics for the giant system, while the reflected and transmitted light retains non-Markovian properties. In this scenario, we show that a nonreciprocal DPT can occur, and the density of reflected photons is influenced by nonlocal correlations. Our findings may inspire new applications and advancements in nonreciprocal photon devices and deterministic photon generation. Furthermore, we present a way to explore nonlocal correlations through photon reflection, offering fundamental insights into phenomena in the DPT .
@@ Acknowledgements.
This work was supported by the National Natural Science Foundation of China under Grant No. U2141237.
Supplemental Materials
In this supplemental material, we show a detailed derivation of the single- and two-photon S matrix, using the method developed in Shi et al. (2015). First, we rewrite the $S$ matrix for reflection and transmission photons in forms of the two modes $A_{\pm k}$ defined in Eq. (11) as
@@
+$+\{U_{p_{2},22}^{\ast}U_{p_{1},12}^{\ast}U_{k_{2},22}U_{k_{1},12}\left\langle 0\right|A_{p_{1}+}e^{-i2HT}A_{k_{1}+}^{\dagger}\left|0\right\rangle e^{i\left(k_{1}+p_{1}\right)T}\delta\left(k_{2}-p_{2}\right)$
@@ I S-matrix in coherent-state presentation
II Single-photon scattering
@@
where the Green function
@@
$\Sigma\left(\omega\right)\equiv\int\mathrm{d}k\frac{\tilde{V}_{k}^{2}}{\omega-k+i0^{+}}=-i2\gamma\left(1+\cos\phi e^{i\left(k_{0}+\omega\right)L}\right).$
Here, the real part Re$\left[\Sigma\left(\omega\right)\right]=2\gamma\cos\phi\sin\left[\left(k_{0}+\omega\right)L\right]$ of the self energy is the frequency correction to the giant system due to the waveguide-emitter coupling, while the imaginary part -Im$\left[\Sigma\left(\omega\right)\right]=2\gamma\left(1+\cos\phi\cos\left[\left(k_{0}+\omega\right)L\right]\right)$ is the waveguide-induced decay rate. We note that in the Markov approximation, $\Sigma\left(\omega\right)$ is taken as the frequency-independent self energy $\Sigma\left(0\right)$, which is generally valid when $\gamma L\ll 1$.
@@
$\left\langle 0\right|a_{Rk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=\delta\left(k_{i}-k_{f}\right)-i2\gamma\left|e^{i\phi}+e^{i\left(k_{0}+k_{i}\right)L}\right|^{2}G\left(k_{i}\right)\delta\left(k_{i}-k_{f}\right)$ (S21)
@@
$\left\langle 0\right|a_{Lk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=-i\gamma\left(e^{-i\phi}+e^{i\left(k_{0}+k_{i}\right)L}\right)\left(e^{i\phi}+e^{i\left(k_{0}+k_{i}\right)L}\right)G\left(k_{i}\right)\delta\left(k_{i}+k_{f}\right).$ (S22)
III Two-photon scattering
@@
[…]
@@
$+i\frac{\gamma^{2}}{2\pi}\left[e^{i\phi}+e^{i\left(k_{0}+k_{1}\right)L}\right]\left[e^{i\phi}+e^{i\left(k_{0}+k_{2}\right)L}\right]TG\left(k_{1}\right)G\left(k_{2}\right)F_{t}\left(X_{1},X_{2}\right)$
@@
$\equiv$ $\frac{1}{\pi}\int\mathrm{d}pe^{ipX_{1}}e^{i\left(k_{1}+k_{2}-p\right)X_{2}}\left(e^{-i\phi}+e^{-i\left(k_{0}+p\right)L}\right)\left(e^{-i\phi}+e^{-i\left(k_{0}+2\bar{k}_{i}-p\right)L}\right)G\left(p\right)G\left(k_{1}+k_{2}-p\right)$
[…]
$=$ $-\frac{ie^{i\bar{k}_{i}\left(X_{1}+X_{2}\right)}}{\bar{k}_{i}+i2\gamma}\sum_{mn}\frac{1}{m!}\left(\frac{-i\gamma\cos\phi}{\bar{k}_{i}+i2\gamma}e^{i\left(k_{0}+\bar{k}_{i}\right)L}\right)^{m+n}\{\left(e^{-i2\phi}+e^{-i2\left(k_{0}+\bar{k}_{i}\right)L}\right)e^{\left(i\bar{k}_{i}-2\gamma\right)X_{mn}}\sum_{l=0}^{n}\tilde{C}_{mnl}\left(k\right)\Theta\left(X_{mn}\right)$

### v2

@@ II.3 Non-Makovian dissipative phase transition
Figure 3: Dissipative phase transition occurs only when the giant system is driven by right-moving input. (a) Photon number (blue solid line, left axis) in the cavity and the second-order correlation function (red line with circles, right axis), as functions of the driving strength (b) Photon number density (blue lines, left axis) of the reflected field and the two-point correlation (red lines with circles, right axis) for different distance $\gamma d=0.63$ (solid lines), $\gamma d=63$ (dashed lines), and $\gamma d=6300$ (dotted lines). Other parameters are $U=\gamma$ and $k_{i}=10\gamma$.
To quantitatively assess the impact of non-Markovian effect on the DPT, we set $k_{0}d=\pi/2$ and $d=2n\pi/k_{i}$ to satisfy the chirality condition in Eq. (3), and examine the DPT via $\rho_{L}$ for various values of $d$. As $d$ increases, the DPT in the reflected field becomes more pronounced, as illustrated in Fig. 3(b). Additionally, a peak emerges near the critical point where the Liouvillian gap reaches a minimum (vanishing only in the limit of infinite photon number), resulting in maximal correlations. $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$. As $d$ further increases, the peak in $\rho_{L}$ approaches the critical point, and the magnitude of the correlator $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$ diminishes because only the eigenstate with the smallest gap contributes significantly. Specifically, when $d$ exceeds the Liouvillian gap significantly, such as $d=2000\pi/k_{i}$, represented by the dotted lines in Fig. 3(b), no correlations are present and $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}$ approaches $\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$. In this scenario, the photon density $\rho_{L}$ reaches its maximum since $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}$ deviates most significantly from $\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}$.
It is also worth noting that the reflected field behaves differently from the cavity field beyond the critical point. While the cavity photon number continues to grow with the driving amplitude $\Omega_{0}$, the reflected photon density $\rho_{L}$decreases after the critical point. This reduction arises from the suppression of the first-order correlator $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$ after the transition. The diminishing correlations are further evidenced by the second-order correlation $g^{\left(2\right)}\left(0\right)$ approaching unity, as shown in Fig. 3(a). After the critical point, $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}\approx\left|\left\langle b\right\rangle_{S}\right|^{2}$ an $\rho_{L}\approx 2\gamma\left(\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}-\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}\right)$. Since both $\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}$ and $\left|\left\langle b\right\rangle_{\mathrm{S}}\right|^{2}$ increase with the driving amplitude $\Omega_{0}$, and the cavity field approaches a coherent state in the large $\Omega_{0}$ limit, the reflected photon density $\rho_{L}$ consequently decreases and eventually vanishes. These findings demonstrate that this system offers a valuable platform for exploring non-Markovian effects in DPTs.
II.4 Connection to experimental realizations
In this section, we present possible experimental implementations and discuss the impact of small deviations from the condition in Eq. (3) on the direction-dependent photon correlations and the nonreciprocal features of the dissipative phase transition.
II.4.1 Implementation
Our theoretical proposal can be implemented in superconducting circuit platforms. As demonstrated in a recent experiment Joshi et al. (2023), a giant system is coupled to a waveguide with a controllable phase difference between two spatially separated coupling points. As shown in Fig. 4(a), two sets of auxiliary resonators $C$ and $R$ are used. $C$ are tunable couplers, whose frequencies are modulated periodically via flux bias lines. $R$ are used to filter undesire emission into the waveguide. The phase difference $\phi_{\mathrm{L}}-\phi_{\mathrm{R}}$ corresponds to the phase $\phi$ in our model. The giant Kerr cavity can be realized as an LC circuit incorporating a nonlinear element like a Josephson junction Blais et al. (2021), realizing a finite Kerr nonlinearity $U$.
Figure 4: Schematics of the implementations of the theoretical model using superconducting circuits. In (a), four auxiliary resonators are employed. The frequencies of two of them—the couplers (CC)—are periodically modulated with a phase difference $\phi=\phi_{L}-\phi_{R}$. In (b), both the cavity frequency and its coupling to the waveguide at one end are periodically modulated. The coupling modulation is realized via a variable coupler composed of inductors and a SQUID.
[…]
$\tilde{H}_{S}=\left(\Delta+A\cos\Delta t\right)b^{{\dagger}}b+Ub^{{\dagger}2}b^{2}$ (20)
is the giant system’s Hamiltoian with modulation amplitude $A$ and
$\tilde{H}_{1}$ $=$ $\int_{-\infty}^{+\infty}\mathrm{d}k[V_{1}\left(a_{\mathrm{L}k}+a_{\mathrm{R}k}\right)b^{{\dagger}}\cos\left(\Delta t+\phi^{\prime}\right)+$ (21)
$V_{2}\left(a_{\mathrm{L}k}e^{-i\left(k_{0}-k\right)L}+a_{\mathrm{R}k}e^{i\left(k_{0}+k\right)L}\right)b^{{\dagger}}+\mathrm{H.c.}$
is its linear coupling to the waveguide, with coupling constants $V_{1}$ and $V_{2}$. In the rotating frame with respect to $\tilde{H}_{0}\left(t\right)=\left(\Delta+A\cos\Delta t\right)b^{{\dagger}}b$, the Hamitltonian becomes
$H=U^{\dagger}\left(t,0\right)\tilde{H}\left(t\right)U\left(t,0\right)-\tilde{H}_{0}\left(t\right)\approx H_{0}+H_{1},$ (22)
where the time evolution operator $U\left(t,0\right)=\mathcal{T}e^{-i\int_{0}^{t}\tilde{H}_{0}\left(t^{\prime}\right)dt^{\prime}}=e^{-i\left(\Delta t+\frac{A}{\Delta}\sin\Delta t\right)b^{{\dagger}}b}$. Here, we have ignored the fast oscillation terms assuming that $\left|V_{1,2}J_{n}\left(\frac{A}{\Delta}\right)\right|^{2}\ll\left|\Delta\right|$, and $V$ and $\phi$ are connected to the modulation parameters as
$V$ $=$ $\frac{V_{1}}{2}\left|J_{2}\left(\frac{A}{\Delta}\right)e^{i\phi^{\prime}}+J_{0}\left(\frac{A}{\Delta}\right)e^{-i\phi^{\prime}}\right|$ (23)
$=$ $V_{2}J_{1}\left(\frac{A}{\Delta}\right)$
and
$\phi=\text{Arg}\left(J_{2}\left(\frac{A}{\Delta}\right)e^{i\phi^{\prime}}+J_{0}\left(\frac{A}{\Delta}\right)e^{-i\phi^{\prime}}\right).$ (24)
II.4.2 Deviations from the chirality condition in (3)
When the condition in Eq. (3) is not fulfilled, the properties of the transmitted photons under right-moving driving are altered, and the behavior of transmitted photons under left-moving driving becomes nontrivial, since the giant system can now be driven by the left-moving input. For a small phase deviation $\Delta\phi$ from the value specified in Eq. (3), the second-order correlation functions $g_{R,L}^{\left(2\right)}\left(0\right)$ for the transmitted photons are shown in Figs. 5(a) and (b). The phase is set to $\phi=\mathrm{\pi}/15+\Delta\phi$, with other parameters the same as in Figs. 2(a) and (b).
Compare to the cases in Figs. 2(a) and (b) with right-moving driving, the transmission probability is slight less than $1$, and the second-order correlation function $g_{R}^{\left(2\right)}\left(0\right)$ shows only minor deviation for small $\Delta\phi=\pi/100$ and $\pi/50$. Meanwhile, $g_{L}^{\left(2\right)}\left(0\right)$ under left-moving driving exhibits nontrivial statistics, with $g_{L}^{\left(2\right)}\left(0\right)$ slightly deviating from 1. Notably, the deviations in $g_{R,L}^{\left(2\right)}\left(0\right)$ diminish for smaller $U$, because only modes with smaller effective decay rate $\gamma_{k}$ contribute to nontrivial second-order correlation functions. For smaller $U$, fewer of these modes are effectively excited, resulting in a reduced sensitivity to $\Delta\phi$. Therefore, for sufficiently small $\Delta\phi$, $g_{L}^{\left(2\right)}\left(0\right)$ for left-moving driving remains approximately 1, and nontrivial photon statistics are generated predominantly for a single incident direction.
Figure 5: Deviations from the chirality condition. Second-order correlation function $g_{R,L}^{(2)}(0)$ of the transmitted photons in the weak-driving limit, and the density of the reflected photons in the strong-driving limit, when the chirality condition is slightly violated. The driving field is right-moving in (a) and (c), and left-moving in (b) and (d). In (a) and (b), the distance is $d=100/\gamma$, with $U=\gamma$ (blue solid and red dashed curves) and $0.1\gamma$ (yellow dotted and purple dash-dotted curves). The phase deviations are $\Delta\phi=\pi/100$ (blue solid and yellow dotted curves) and $\pi/50$ (red dashed and purple dash-dotted curves). The inset in (b) shows the transmission probability, with the legend consistent with that in (a) and (b). In (c) and (d), the phase is fixed at $\phi=\pi/2$, the distance is $d=6300/\gamma$, and different deviations of the propagation phase are considered. Other parameters are the same as in Fig. 3.
We now consider the impact of a small deviation from Eq. (3) on the DPT under the strong-diving limit. To ensure Markovian dynamics in the giant system, $\phi$ is set to be $\mathrm{\pi}/2$, and a small deviation $\Delta\theta$ is introduced, making the propagation phase $\theta_{\mathrm{R}k_{i}}=\mathrm{\pi}/2+\Delta\theta$. With other parameters identical to those in Fig. 3, the density of reflected photons under right-moving driving shows negligible change, as illustrated in Fig. 5(c). Under left-moving driving, the reflected photon density $\rho_{R}$ also exhibits a DPT, but at a critical pumping amplitude $\Omega_{0}$ that is more than 10 times larger, as shown in Fig. 5(d). This is because the effective driving amplitude $\Omega_{L}\left(k_{i}\right)\ll\Omega_{R}\left(k_{i}\right)$. As a result, if we focus on the critical behavior within the driving amplitude range around the critical point for right-moving driving, the DPT remains nonreciprocal.
III Conclusion
In summary, we have explored the non-Markovian multiphoton dynamics of a driven giant system under chiral coupling conditions, achieved by matching the phase difference $\phi$ between coupling points to the propagation phase of the input field. Focusing on a nonlinear giant cavity coupled unidirectionally to a waveguide, we demonstrated that in the weak-driving (few-photon) regime, strongly nonclassical photon statistics can be deterministically generated—even in the presence of strong dissipation and weak nonlinearity. This arises from coherent interference between forward- and backward-propagating components within the nonlocal coupling region and is tunable via the coupling-point separation, reflecting the intrinsic non-Markovianity of the system. At the special point $\phi=\pm\mathrm{\pi}/2$, interference between coupling points vanishes, rendering the system’s internal dynamics exactly Markovian. Nonetheless, the output fields remain non-Markovian due to the extended coupling structure. In this regime, we uncover a nonreciprocal dissipative phase transition in the reflected field, governed by nonlocal photon correlations. The physical implementations are presented, along with a discussion of the impact of small deviations from the chirality condition. Our study can be extended to casese with $N$ coupling legs (see Supplementary Note 4). The findings offer a robust platform for engineering chiral, non-Markovian multiphoton states in driven-dissipative settings. This framework is broadly applicable to a variety of nonlocally coupled quantum systems, including giant cavities interfaced with atomic ensembles and central spin models coupled to spin chains, providing a foundation for scalable quantum state control in open quantum networks.
IV Methods
IV.1 Numerical simulations
The dissipative phase transitions shown in Figs. 3 and 5 are obtained by solving the master equation (15) and using the input-output formalism (18). the master equation (15) is solved in the superspace, where the density matrix $\rho$ is reshaped as a column vector $\vec{\rho}$ of dimension $D^{2}$. Here, $D$ is the Hilber space dimension, with D1D-1 corresponding to the cutoff of the boson number for mode $b$. In the superspace, the Liouvillian $\mathcal{L}$, defined as
$\mathcal{L}\rho\equiv-i\left[H_{S},\rho\right]+2\gamma\left(2b\rho b^{\dagger}-\left\{b^{\dagger}b,\rho\right\}\right),$ (25)
is represented as a $D^{2}\times D^{2}$ matrix
$\mathcal{L}_{M}$ $=$ $-i\left(H_{S}\otimes I_{D}-I_{D}\otimes\left(H_{S}\right)^{T}\right)+$ (26)
$2\gamma\left(2b\otimes\left(b^{\dagger}\right)^{T}-b^{\dagger}b\otimes I_{D}-I_{D}\otimes b^{\dagger}b\right),$
where $I_{D}$ is the $D$-dimensional identity matrix, and $\left(\bullet\right)^{T}$ denotes the matrix transpose.
The steady state solution $\vec{\rho}_{ss}$ of the master equation can be obtained by diagonalizing the Liouvillian matrix $\mathcal{L}_{M}$ and identifying the zero-eigenvalue state. The steady state must satisfy the normalization condition $u_{0}\vec{\rho}_{ss}=1$, which follows from Tr[$\left[\rho\right]=1$, where $u_{0}$ is generated by reshaping the $D$-dimensional identity matrix $I_{D}$ into a $1\times D^{2}$ row vector. Alternatively, $\vec{\rho}_{ss}$ can be found from the equation $\mathcal{L}_{M}\vec{\rho}_{ss}=0$. Note that $\mathcal{L}_{M}$ is not reversible (its rank is less than $D^{2})$, and that is why it has zero eigenvalues. Physically, this reflects Tr[$\left[\rho\right]=1$, which implies Tr[$\left[\mathcal{L}\rho\right]=0$. To incorporate the normalization condition Tr[$\left[\rho\right]=1$, one replaces the row corresponding to the diagonal-element of $\vec{\rho}$ ($1,D+1,2D+2,...,D^{2}$)with $u_{0}$. Without loss of generality, we replace the last row of $\mathcal{L}_{M}$ with $u_{0},$ forming a modified matrix $\mathcal{L}_{M}^{\prime}$. The steady state is then given by $\vec{\rho}_{ss}=\left(\mathcal{L}_{M}^{\prime}\right)^{-1}v_{0}$, where $v_{0}$ is a $D^{2}\times 1$ column vector with $1$ in the last element and zeros elsewhere.
Once the steady state $\vec{\rho}_{ss}$ is obtained, the long-term-average of any system operator can be calculated. For example, the mean value photon number value is given by
$\langle b^{{\dagger}}b\rangle_{S}=u_{0}\left(b^{{\dagger}}b\otimes I_{D}\right)\vec{\rho}_{ss}.$ (27)
Multi-time correlations function can be evaluated using the quantum regression theorem Gardiner and Zoller (2004); Walls and Milburn (2008). For instance, the two-time correlation function in Eq. (19) can be computed as
$\left\langle b^{{\dagger}}\left(d\right)b\left(0\right)\right\rangle=u_{0}\left(b^{{\dagger}}\otimes I_{D}\right)e^{\mathcal{L}_{M}d}\left(b\otimes I_{D}\right)\vec{\rho}_{ss}.$ (28)
V Data availability
Data that support the findings of this study are available from the corresponding author upon reasonable request.
VI Code availability
The codes used in this study are available from the corresponding author upon reasonable request.
VII Acknowledgements
@@ Acknowledgements.
The author would like to thank Carlos Navarrete-Benlloch for fruitful discussions on the physical implementation.
Supplemental information for “Non-Markovian Multiphoton Chiral Dynamics with Giant Systems ”
Supplementary Note1: S-matrix in coherent-state presentation
We first rewrite the $S$ matrix for reflection and transmission photons in forms of the two modes $A_{\pm k}$ defined in Eq. (11) as
@@
$+\{U_{p_{2},22}^{\ast}U_{p_{1},12}^{\ast}U_{k_{2},22}U_{k_{1},12}\left\langle 0\right|A_{p_{1}+}e^{-i2HT}A_{k_{1}+}^{\dagger}\left|0\right\rangle e^{i\left(k_{1}+p_{1}\right)T}\delta\left(k_{2}-p_{2}\right)$
@@
Supplementary Note 2: Single-photon scattering
@@
where the Green’s function
@@
$\Sigma\left(\omega\right)\equiv\int\mathrm{d}k\frac{\tilde{V}_{k}^{2}}{\omega-k+i0^{+}}=-i2\gamma\left(1+\cos\phi e^{i\left(k_{0}+\omega\right)d}\right).$ (S21)
[…]
@@
$\left\langle 0\right|a_{Rk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=\delta\left(k_{i}-k_{f}\right)-i2\gamma\left|e^{i\phi}+e^{i\left(k_{0}+k_{i}\right)d}\right|^{2}G\left(k_{i}\right)\delta\left(k_{i}-k_{f}\right)$ (S22)
@@
$\left\langle 0\right|a_{Lk_{f}}Sa_{Rk_{i}}^{\dagger}\left|0\right\rangle=-i\gamma\left(e^{-i\phi}+e^{i\left(k_{0}+k_{i}\right)L}\right)\left(e^{i\phi}+e^{i\left(k_{0}+k_{i}\right)d}\right)G\left(k_{i}\right)\delta\left(k_{i}+k_{f}\right).$ (S23)
Supplementary Note 3: Two-photon scattering
@@
$=$ $\int d\omega\sum_{mn}\frac{e^{im\left(k_{0}+\omega\right)L}\left(-i2\gamma\cos\phi\right)^{m}}{\left(\omega+i2\gamma\right)^{m+1}}\frac{e^{in\left(k_{0}+k_{1}+k_{2}-\omega\right)L}\left(-i2\gamma\cos\phi\right)^{n}}{\left(k_{1}+k_{2}-\omega+i2\gamma\right)^{n+1}}$
@@
$+i\frac{\gamma^{2}}{2\pi}\left[e^{i\phi}+e^{i\left(k_{0}+k_{1}\right)d}\right]\left[e^{i\phi}+e^{i\left(k_{0}+k_{2}\right)d}\right]TG\left(k_{1}\right)G\left(k_{2}\right)F_{t}\left(X_{1},X_{2}\right)$
@@
$\equiv$ $\frac{1}{\pi}\int\mathrm{d}pe^{ipX_{1}}e^{i\left(k_{1}+k_{2}-p\right)X_{2}}\left(e^{-i\phi}+e^{-i\left(k_{0}+p\right)d}\right)\left(e^{-i\phi}+e^{-i\left(k_{0}+2\bar{k}_{i}-p\right)d}\right)G\left(p\right)G\left(k_{1}+k_{2}-p\right)$
$=$ $\frac{1}{\pi}e^{i2\bar{k}_{i}X_{2}}\sum_{mn}\left(-i2\gamma\cos\phi\right)^{m+n}\left(-\right)^{n+1}e^{imk_{0}L}e^{in\left(k_{0}+2\bar{k}_{i}\right)L}\int dp\frac{e^{ipX_{mn}}\left(e^{-i\phi}+e^{-i\left(k_{0}+p\right)d}\right)\left(e^{-i\phi}+e^{-i\left(k_{0}+2\bar{k}_{i}-p\right)d}\right)}{\left(p+i2\gamma\right)^{m+1}\left(p-2\bar{k}_{i}-i2\gamma\right)^{n+1}}$
$=$ $-\frac{ie^{i\bar{k}_{i}\left(X_{1}+X_{2}\right)}}{\bar{k}_{i}+i2\gamma}\sum_{mn}\frac{1}{m!}\left(\frac{-i\gamma\cos\phi}{\bar{k}_{i}+i2\gamma}e^{i\left(k_{0}+\bar{k}_{i}\right)d}\right)^{m+n}\{\left(e^{-i2\phi}+e^{-i2\left(k_{0}+\bar{k}_{i}\right)d}\right)e^{\left(i\bar{k}_{i}-2\gamma\right)X_{mn}}\sum_{l=0}^{n}\tilde{C}_{mnl}\left(k\right)\Theta\left(X_{mn}\right)$
@@
Supplementary Note 4: $N$ coupling points
With $N$ coupling points with the same coupling strength $V$ and diffirent phase $\phi_{n}$, the interaction between the giant cavity and the waveguide is $H_{1}=\int_{-\infty}^{+\infty}\tilde{V}_{k}\left(b^{{\dagger}}A_{+k}+\mathrm{H.c.}\right)\mathrm{d}k\mathrm{,}$where
$\tilde{V}_{k}=V\sqrt{2\sum_{m,n=0}^{N}e^{i\left(\phi_{m}-\phi_{n}\right)}\cos\left[\left(m-n\right)\left(k_{0}+k\right)d\right]}$ (S38)
and the coupled mode $A_{+k}$ and free mode $A_{-k}$ are defined as
$\left(\begin{array}[]{c}A_{+k}\\ A_{-k}\end{array}\right)=U_{k}\left(\begin{array}[]{c}a_{L-k}\\ a_{Rk}\end{array}\right),$ (S39)
with the unitary transformation
$U_{k}=\frac{V}{\tilde{V}_{k}}\left(\begin{array}[]{cc}\sum_{n=1}^{N}e^{i\phi_{n}}e^{-i\left(n-1\right)\left(k_{0}+k\right)d}&\sum_{n=1}^{N}e^{i\phi_{n}}e^{i\left(n-1\right)\left(k_{0}+k\right)d}\\ \sum_{n=1}^{N}e^{-i\phi_{n}}e^{-i\left(n-1\right)\left(k_{0}+k\right)d}&-\sum_{n=1}^{N}e^{-i\phi_{n}}e^{i\left(n-1\right)\left(k_{0}+k\right)d}\end{array}\right).$ (S40)
The self-energy is acquired as
$\Sigma\left(\omega\right)=-i\gamma\left[N+2\sum_{m>n}^{N}\cos\left(\phi_{m}-\phi_{n}\right)e^{i\left(m-n\right)\left(k_{0}+\omega\right)d}\right]$ (S41)
where $\gamma=2\pi V^{2}$. For photon scattering, one only needs to replace the self-energy in the $N=2$ case by $\Sigma\left(\omega\right)$ in Eq. (S41), as well as changing $\eta_{Rk}$ by
$\eta_{Rk}=\sum_{n=1}^{N}e^{i\phi_{n}}e^{i\left(n-1\right)\left(k_{0}+k\right)d}.$ (S42)
For example, the transmission amplitude $t\left(k_{i}\right)=\left\langle k_{f}^{R}\right|S\left|k_{i}^{R}\right\rangle$ is $t\left(k_{i}\right)=\delta\left(k_{i}-k_{f}\right)-i\gamma\left|\eta_{Rki}\right|^{2}G\left(k_{i}\right)\delta\left(k_{i}-k_{f}\right)$, where the Green’s function $G\left(\omega\right)=\frac{1}{\omega-\Sigma\left(\omega\right)}$. The S-matrix for two photons with right-moving incident momenta $k_{1}$and $k_{2}$, and right moving output momenta $p_{1}$ and $p_{2}$ is
$S_{p_{2}p_{1},k_{2}k_{1}}$ $=$ $\left[\delta\left(k_{1}-p_{1}\right)\delta\left(k_{2}-p_{2}\right)+\delta\left(k_{1}-p_{2}\right)\delta\left(k_{2}-p_{1}\right)\right]t\left(k_{1}\right)t\left(k_{2}\right)$ (S43)
$-i\frac{\gamma^{2}}{\pi}\eta_{Rp_{2}}^{\ast}\eta_{Rp_{1}}^{\ast}\eta_{Rk_{2}}\eta_{Rk_{1}}G\left(k_{1}\right)G\left(k_{2}\right)T_{S}G\left(p_{1}\right)G\left(p_{2}\right)\delta\left(k_{1}+k_{2}-p_{1}-p_{2}\right),$
where the $T_{s}$-matrix is
The condition where the dynamics of the giant system is Markovian is equivalent to the condition when $\tilde{V}_{k}$ independent of $k$. It gives
$\phi_{1}=\pm\frac{\pi}{2}$ (S45)
and
$\sum_{n=1}^{N-1}\cos\left(\phi_{n}-\phi_{n+1}\right)=0,\sum_{n=1}^{N-2}\cos\left(\phi_{n}-\phi_{n+2}\right)=0,\sum_{n=1}^{N-3}\cos\left(\phi_{n}-\phi_{n+3}\right)=0,$... (S46)
For simplicity, we have set $\phi_{N}=0$. When these conditions can be fulfilled, the dynamics of the giant system can be exactlydepicted by the master equation
$\partial_{t}\rho=-i\left[H_{S},\rho\right]+N\gamma\left(2b\rho b^{\dagger}-\left\{b^{\dagger}b,\rho\right\}\right),$ (S47)
where $\gamma=2\pi V^{2}$. The input-output formalism read
$a_{out}^{l}\left(t\right)=a_{in}^{l}\left(t\right)-i\sqrt{\gamma}\sum_{n=1}^{N}e^{-i\phi_{n}}e^{i\left(n-1\right)k_{0}L}b\left(t-\left(n-1\right)d\right),$ (S48)
$a_{out}^{r}\left(t\right)=a_{in}^{r}\left(t\right)-i\sqrt{\gamma}\sum_{n=1}^{N}e^{-i\phi_{n}}e^{-i\left(n-1\right)k_{0}L}b\left(t+\left(n-1\right)d\right).$ (S49)
With $M=3$, the conditions can be fulfilled with $\phi_{1}=\pm\pi/2$, $\phi_{2}=\mp\pi/4$, resulting in $\tilde{V}_{k}=\sqrt{6}V$. In this case, the master equation reads
$\partial_{t}\rho=-i\left[H_{S},\rho\right]+3\gamma\left(2b\rho b^{\dagger}-\left\{b^{\dagger}b,\rho\right\}\right),$ (S50)
with the input-output formalism as
$a_{out}^{l}\left(t\right)=a_{in}^{l}\left(t\right)-i\sqrt{\gamma}\left[e^{-i\phi_{1}}b\left(t\right)+e^{-i\phi_{2}}e^{ik_{0}d}b\left(t-d\right)+e^{i2k_{0}L}b\left(t-2d\right)\right],$ (S51)
$a_{out}^{r}\left(t\right)=a_{in}^{r}\left(t\right)-i\sqrt{\gamma}\left[e^{-i\phi_{1}}b\left(t\right)+e^{-i\phi_{2}}e^{-ik_{0}d}b\left(t+d\right)+e^{-i2k_{0}L}b\left(t+2d\right)\right].$ (S52)
We note that for $N>2$, the chiral condition is not as simple as in the $N=2$. In fact, for a given incident momentum $k_{i}$, one need to solve a transcendental equation $\sum_{n=1}^{N}e^{i\phi_{n}}e^{-i\left(n-1\right)\left(k_{0}-k\right)d}=0$ to determine the phase $\phi_{n}$.